Determine the volume occupied by 4.26 g of CO2 gas at STP. Use the equation: V= nRTP



1
Expert's answer
2022-03-22T16:27:05-0400

Solution:

STP = Standard Temperature and Pressure (T = 273.15 K and P = 1 atm)

R = universal gas constant = 0.0821 L atm K−1 mol−1


The molar mass of CO2 is 44.01 g/mol

Therefore,

Moles of CO2 = (4.26 g CO2) × (1 mol CO2 / 44.01 g CO2) = 0.0968 mol CO2

n(CO2) = 0.0968 mol


Ideal gas law can be used:

PV = nRT

To find the volume, solve the equation for V:

V = nRT / P

V = (0.0968 mol × 0.0821 L atm K−1 mol−1 × 273.15 K) / (1 atm) = 2.17 L

V = 2.17 L


Answer: The volume of CO2 is 2.17 L